# Shaocheng Liu and Qi Chen

### v2

@@ Abstract
Characterization of entropy functions is of fundamental importance in information theory. By imposing constraints on their Shannon outer bound, i.e., the polymatroidal region, one obtains the faces of the region and entropy functions on them with special structures. In this series of two papers, we characterize entropy functions on the $2$-dimensional faces of the polymatroidal region of degree $4$. In Part I, we formulate the problem, enumerate all $59$ types of $2$-dimensional faces of the region by an algorithm, and fully characterize entropy functions on $49$ types of them. Among them, those non-trivial cases are mainly characterized by the graph-coloring technique. The entropy functions on the remaining $10$ types of faces will be characterized in Part II, among which $8$ types are fully characterized, and $2$ types are partially characterized.
@@ I Introduction
The region in $\mathcal{H}_{n}$ bounded by such inequalities, denoted by $\Gamma_{n}$, is called the polymatroidal region of degree $n$. Thus, $\Gamma_{n}$ is an outer bound on $\Gamma^{*}_{n}$. For more about entropy functions, we referred the readers to [1, Chapter 13-15].
@@
Traditionally, entropy functions are characterized by information inequalities. Those inequalities derived by polymatroidal axioms are called Shannon-type, as they correspond to the non-negativity of Shannon information measures. Since 1998, a series of non-Shannon-type information inequalities, among which Zhang-Yeung inequality is the first one [2], were discovered [3][4][5]. Thus $\overline{\Gamma^{*}_{n}}$, the closure of $\Gamma^{*}_{n}$, is strictly included in $\Gamma_{n}$ when $n\geq 4$. Each information inequality determines an outer bound of $\Gamma^{*}_{n}$, as those set functions in $\mathcal{H}_{n}$ dissatisfy it must be located outside $\Gamma^{*}_{n}$. In this series of two papers, we develop a system of entropy function characterization from the perspective of faces of $\Gamma_{n}$, which covers the traditional inequality characterization.
@@
By definition, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$. Thus, each Shannon-type information inequality determines a face $F$ of $\Gamma_{n}$. It is natural to characterize entropy functions on the specific $F$ of $\Gamma_{n}$ (See Subsection II-B for details on the faces of a polyhedral cone). Let $F^{*}\triangleq F\cap\Gamma^{*}_{n}$ be the set of all entropy functions in $F$. In the following, to determine the entropy functions on $F$, or the region $F^{*}$, we will call it characterize $F$ for short. A non-Shannon-type information inequality can be considered as an outer bound on $F^{*}$ when $F=\Gamma_{n}$ itself, the improper face of $\Gamma_{n}$. A constrained non-Shannon-type information inequality is an outer bound on $F^{*}$ when $F$ is the face determined by the constraints that are equalities obtained by setting some Shannon-type inequalities as equalities. When $F$ is an extreme ray, i.e., a $1$-dimensional face of $\Gamma_{n}$, if it contains a matroid, entropy functions on $F$ are called matroidal entropy functions, and they can be fully characterized by the probabilistically characteristic set of the matroid [6][7][8]. Matúš fully characterized the first non-trivial 2-dimensional face of $\Gamma_{n}$ in 2006 [9]. It is a $2$-dimensional face of $\Gamma_{3}$. In 2012, Chen and Yeung characterized another $2$-dimensional face of $\Gamma_{3}$ [10]. They are the only two types of non-trivial faces of $\Gamma_{3}$ that need to be characterized. To the best of our knowledge, so far, there is no fully characterized no-trivial $3$-dimensional faces. However, partial characterizations of $3$-dimensional faces of $\Gamma_{3}$ can be found in [11][12].
@@
Many information-theoretic problems can be considered as applications of entropy function characterizations on faces of $\Gamma_{n}$. In a series of three papers [13, 14, 15], Matúš and Studneý solved the probabilistic conditional independence problem for four random variables. Note each class of conditional independence constraints, which is called a semimatroid in their papers, determines a face $F$ of $\Gamma_{n}$. The solution to this problem, i.e., the probabilistic-representability of a semimatroid determines whether the relative interior of the corresponding face $F$ interset with $\Gamma^{*}_{n}$. This problem thus can be considered as partial characterizations of the faces of $\Gamma_{n}$. In [16], Yan, Yeung, and Zhang proved a formula involving $\Gamma_{n}^{*}$ for the capacity of multi-source multi-sink network coding. Those constraints in the formula induced by network topology and source independency form a face of $\Gamma_{n}$, which shows that this holy-grail network coding problem corresponds to the entropy function characterizations on this face. For the secret-sharing problem, see [17] for example, the perfect secrecy and decoding correctness conditions of an access structure determine a face $F$ on $\Gamma_{n}$, and the information ratio of the secret-sharing problem can be considered as an optimization problem whose feasible region is $F\cap\Gamma^{*}_{n}$. Other problems, such as distributed data storage [18], coded caching [19], and Markov random fields [20] are also related to the entropy regions on the faces of $\Gamma_{n}$.
In this series of two papers, we characterize entropy functions on the $2$-dimensional faces of $\Gamma_{4}$. In Part I, we enumerate all $59$ types of $2$-dimensional faces of $\Gamma_{4}$ by an algorithm and completely characterize $49$ types of them. Among them, $13$ types are all-entropic faces, while $7$ are non-entropic ones. The characterization of $17$ types can be obtained by the $2$ types of non-trivial faces of $\Gamma_{3}$ characterized in [9] and [10]. Of the $22$ types of non-trivial faces, $12$ will be characterized in Part I.
@@
The primary technique we adopt to characterize these non-trivial faces is the graph-coloring in [9]. For a $3$-dimensional random vector $(X_{i},i\in N_{3})$, if $X_{1}$ and $X_{2}$ are independent, and $X_{3}$ is a function of $(X_{1},X_{2})$, $X_{1}$ and $X_{2}$ can be considered distributed on the two sets of vertices $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ of a bipartite graph with edges those $(x_{1},x_{2})$ such that $p(x_{1},x_{2})>0$, and these edges are colored by $x_{3}\in\mathcal{X}_{3}$. In our work, when the number of the random variables becomes four, we color bipartite graphs by two color systems or color a tripartite graph by one color system.
@@
The remaining $10$ cases of faces will be characterized in Part II [21] of the two-paper series. Among them, $8$ faces are fully and $2$ are partially characterized. The characterizations breed new combinatorial-design structures, generalizing $\rm VOA$s introduced in [6] [7].
@@ II-B Faces of a polyhedral cone
each of which determines a facet of $\Gamma_{n}$ [1, Chapter 14]. When $n=4$, it can be checked that there are $28$ elemental inequalities, or $28$ facets of $\Gamma_{4}$.
@@
where $\mathbf{r}$ is the minimal integer polymatroid in the ray. Therefore, in this paper, we refer to the minimal integer polymatroid $\mathbf{r}$ as the extreme ray $E_{\mathbf{r}}$ containing it unless otherwise specified. Note that some of these integer polymatroids are matroids. The $41$ extreme rays can be classified into the following $11$ types.
@@
for $\mathcal{W}^{\alpha}_{2}$ with $\alpha\subseteq N_{4}$, $|\alpha|=2$, it is called a wheel matroid with order $2$, [^1] We adopt the notation and terminology in [22, Section 8.4] and it is a matroid with two parallel elements in $\alpha$, and each element in $\alpha$ and the other two elements in $N_{4}$ form a $U_{2,3}$;
@@
for $\hat{U}^{i}_{2,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{2,5}$, [^2] See [25, Theorem 1.3.6] or [26, Theorem 4] for the definition of free expansion. and it’s rank function is defined by
@@
for $\hat{U}^{i}_{3,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{3,5}$, and it’s rank function is defined by
@@
for $V^{\alpha}_{8}$ with $\alpha\subseteq N_{4}$ and $|\alpha|=2$, it is a polymatroid whose free expansion is the Vámos matroid, and it’s rank function is defined by
@@
It can be seen that for an extreme ray in the form $E^{\alpha}$ with $\alpha\subseteq N_{4}$, it is in a type with $\binom{4}{|\alpha|}$ extreme rays, and each extreme ray in the type can be obtained from each other by permuting the indices in $N_{4}$. For a specific extreme ray, say $U^{\{1,2\}}_{1,2}$, for simplicity, we will drop the bracket and comma of the set in the superscript, and write it as $U^{12}_{1,2}$.
@@ II-C Entropy functions on the extreme rays of $\Gamma_{4}$
In the following, extreme rays are written in the form in (II.B.3). For each type of extreme rays , we only consider one of them. Throughout this paper, for any random vector $(X_{i}:i\in N_{4})$, each $X_{i}$ is assumed distributed on a finite set $\mathcal{X}_{i},i\in N_{4}$. For any marginal distribution of $(X_{i}:i\in N_{4})$, say $p_{X_{2}X_{3}}(x_{2},x_{3})$ with $(x_{2},x_{3})\in\mathcal{X}_{2}\times\mathcal{X}_{3}$, we usually drop its subscription and write it as $p(x_{2},x_{3})$ if there is no ambiguity. We may further simplify it as $p(x_{12})$.
@@
For $E=U^{123}_{2,3},\mathcal{W}^{14}_{2},U_{3,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log k$ for integer $k\geq 1$.
@@
([6, Proposition 2]) For $E=U_{2,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log k$ for positive integer $k\neq 2,6$.
@@
Extreme rays $E$ Theorems Figures of entropy region $E^{*}=E\cap\Gamma^{*}_{4}$ Description
@@
$U^{\alpha}_{2,3}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log k,\ k>0,k\mathbb{Z}\}$
$\mathcal{W}^{\alpha}_{2}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log k,\ k>0,k\in\mathbb{Z}\}$
$U_{2,4}$ II-C.3 [Uncaptioned image] $\{a\mathbf{r}:a=\log k,k>0,k\in\mathbb{Z},k\neq 2,6\}$
$U_{3,4}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log k,\ k>0,k\in\mathbb{Z}\}$
$\hat{U}^{i}_{2,5}$ II-C.4 [Uncaptioned image] $\{a\mathbf{r}:a=\log k,\ k>0,k\in\mathbb{Z}\}$
$\hat{U}^{i}_{3,5}$ II-C.5 [Uncaptioned image] $\{a\mathbf{r}:a=\log k,\ k>0,k\in\mathbb{Z}\}$
@@
It is well known that the rays containing a minimal integer polymatroid with rank-$1$ are all-entropic. It has been proven that the a vector $a\in\mathbf{r}$ in $U^{\alpha}_{2,3}$ are entropic if and only if $a=\log{k}$ for $k\in\mathbb{Z}$ [1, Theorem 15.2]. In the ray $U_{2,4}$, $a$ can take the values $\log{k}$ except for $\log 2$ and $\log 6$ [7, Proposition 6]. In the ray $U_{3,4}$, the region is the same as in $U^{\alpha}_{2,3}$. These three cases correspond to the Latin Square, Orthogonal Two Latin Squares, and Latin Cube, respectively. As for the ray $\hat{U}^{i}_{2,5}$ and $\hat{U}^{i}_{3,5}$, they are implied by the following theorems. For the ray $V^{\alpha}_{8}$, the fact that it disobeys the Zhang-Yeung inequality indicates that it lies outside of $\overline{\Gamma_{4}^{*}}$ and, consequently, is non-entropic.
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
@@
For $\mathbf{h}=a\mathbf{r}$, where $\mathbf{r}$ is the rank function of $\hat{U}^{1}_{2,5}$, if we restrict it on $\{2,3,4\}$, we will see that its restriction $\mathbf{h}^{\prime}=a\mathbf{r}_{1}$, where $\mathbf{r}_{1}$ is the rank function of $U_{2,3}$, is entropic only if $a=\log k$ for integer $k>0$.
To prove the “if” part, let $X_{2}$ and $X_{3}$ be independent and uniformly distributed on $\{0,\dots,k-1\}$, and $X_{4}=X_{2}+X_{3}\mod k$. Let $X_{1}=k\cdot X_{2}+X_{3}$. Then $X_{1}$ is uniformly distributed on $\{0,\dots,k^{2}-1\}$ and it can be checked that $(X_{i},i\in N_{4})$ such constructed characterizes $\mathbf{h}=\log{k}\cdot\mathbf{r}$. ∎
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{i,j,k}),\{i,j,k\}=\{2,3,4\}$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),|K|=2,K\nsubseteq\{2,3,4\}.$
Assume that each $X_{i}$ is distributed on $\mathcal{X}_{i}$ and for $x_{i}\in\mathcal{X}_{i}$, $p(x_{i})>0$. For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
By Lemma 1, $X_{2},X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{2},\mathcal{X}_{3}$ and $\mathcal{X}_{4}$ and $H(X_{2})=H(X_{3})=H(X_{4})=\log k$, where $k=|\mathcal{X}_{i}|,i\in\{2,3,4\}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have $H(X_{2})=H(X_{3})=H(X_{4})=a$. Thus $a$ can only take the value $\log k$.
Now we prove that all of these polymatroids are entropic. For $a=\log k$, let $X_{2},X_{3}$ and $X_{4}$ be mutually independent and uniformly distributed on $\{0,1,\dots,k-1\}$. Let
$X_{1}=k[(X_{3}-X_{2})\mod k]+[(X_{4}-X_{2})\mod k].$ (II.C.23)
It follows to check the entropy function of such constructed $(X_{i},i\in N_{4})$ is $a\mathbf{r}$. Given that $X_{2}$, $X_{3}$ and $X_{4}$ are mutually independent and uniformly distributed, we have
$H(X_{2})$ $=H(X_{3})=H(X_{4})=\log k$ (II.C.24)
$H(X_{234})$ $=H(X_{2})+H(X_{3})+H(X_{4})=3\log k.$ (II.C.25)
Then by (II.C.23), we obtain $H(X_{1234})=H(X_{234})$. Dividing both sides of (II.C.23) by $k$ and then rounding them down, we obtain
$\lfloor\dfrac{X_{1}}{k}\rfloor=(X_{3}-X_{2})\mod k$ (II.C.26)

### v3

# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part I: Problem Formulation and More

@@ Abstract
Characterization of entropy functions is of fundamental importance in information theory. By imposing constraints on their Shannon outer bound, i.e., the polymatroidal region, one obtains the faces of the region and entropy functions on them with special structures. In this series of two papers, we characterize entropy functions on the $2$-dimensional faces of the polymatroidal region of degree $4$. In Part I, we formulate the problem, enumerate all $59$ types of $2$-dimensional faces of the region by an algorithm, and fully characterize entropy functions on $49$ types of them. The entropy functions on the remaining $10$ types of faces will be characterized in Part II, among which $8$ types are fully characterized, and $2$ types are partially characterized.
Index terms - entropy function, entropy region, polymatroid, poymatroidal region
@@ I Introduction
The region in $\mathcal{H}_{n}$ bounded by such inequalities, denoted by $\Gamma_{n}$, is called the polymatroidal region of degree $n$. Thus, $\Gamma_{n}$ is an outer bound on $\Gamma^{*}_{n}$. For more about entropy functions, we referred the readers to [2, Chapter 13-15].
@@
Traditionally, entropy functions are characterized by information inequalities. Those inequalities derived by polymatroidal axioms are called Shannon-type, as they correspond to the non-negativity of Shannon information measures. Since 1998, a series of non-Shannon-type information inequalities, among which Zhang-Yeung inequality is the first one [3], were discovered [4, 5, 6]. Thus $\overline{\Gamma^{*}_{n}}$, the closure of $\Gamma^{*}_{n}$, is strictly included in $\Gamma_{n}$ when $n\geq 4$. Each information inequality determines an outer bound on $\Gamma^{*}_{n}$, as those set functions in $\mathcal{H}_{n}$ dissatisfy it must be located outside $\Gamma^{*}_{n}$. In this series of two papers, we develop a system of entropy function characterization from the perspective of faces of $\Gamma_{n}$, which covers the traditional inequality characterization.
@@
By definition, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$. Thus, each Shannon-type information inequality determines a face $F$ of $\Gamma_{n}$. It is natural to characterize entropy functions on the specific $F$ of $\Gamma_{n}$ (See Subsection II-B for details on the faces of a polyhedral cone). Let $F^{*}\triangleq F\cap\Gamma^{*}_{n}$ be the set of all entropy functions in $F$. In what follows, when we refer to determining the entropy functions on $F$ , or the region $F^{*}$, we will simply call this characterizing $F$. A non-Shannon-type information inequality can be considered as an outer bound on $F^{*}$ when $F=\Gamma_{n}$ itself, the improper face of $\Gamma_{n}$. A constrained non-Shannon-type information inequality is an outer bound on $F^{*}$ when $F$ is the face determined by the constraints that are equalities obtained by setting some Shannon-type inequalities as equalities. When $F$ is an extreme ray, i.e., a $1$-dimensional face of $\Gamma_{n}$, if it contains a matroid, entropy functions on $F$ are called matroidal entropy functions, and they can be fully characterized by the probabilistically characteristic set of the matroid [7, 8, 9]. Matúš fully characterized the first non-trivial 2-dimensional face of $\Gamma_{n}$ in 2006 [10]. It is a $2$-dimensional face of $\Gamma_{3}$. In 2012, Chen and Yeung characterized another $2$-dimensional face of $\Gamma_{3}$ [11]. They are the only two types of non-trivial $2$-dimensional faces of $\Gamma_{3}$ that need to be characterized. To the best of our knowledge, so far, there is no fully characterized no-trivial $3$-dimensional faces. However, partial characterizations of $3$-dimensional faces of $\Gamma_{3}$ can be found in [12, 13].
@@
Many information-theoretic problems can be considered as applications of entropy function characterizations on faces of $\Gamma_{n}$. In a series of three papers [14, 15, 16], Matúš and Studneý solved the probabilistic conditional independence problem for four random variables. Note each class of conditional independence constraints, which is called a semimatroid in their papers, determines a face $F$ of $\Gamma_{n}$. The solution to this problem, i.e., the probabilistic-representability of a semimatroid determines whether the relative interior of the corresponding face $F$ intersets with $\Gamma^{*}_{n}$. This problem thus can be considered as partial characterizations of the faces of $\Gamma_{n}$. In [17], Yan, Yeung, and Zhang proved a formula involving $\Gamma_{n}^{*}$ for the capacity of multi-source multi-sink network coding. Those constraints in the formula induced by network topology and source independency form a face of $\Gamma_{n}$, which shows that this holy-grail network coding problem corresponds to the entropy function characterizations on this face. For the secret-sharing problem, see [18] for example, the perfect secrecy and decoding correctness conditions of an access structure determine a face $F$ on $\Gamma_{n}$, and the information ratio of the secret-sharing problem can be considered as an optimization problem whose feasible region is $F\cap\Gamma^{*}_{n}$. Other problems, such as distributed data storage [19], coded caching [20], Markov random fields [21], and relational database [22] are also related to the entropy regions on the faces of $\Gamma_{n}$. Recently, there are a series of works on how to compute the faces of $\Gamma_{n}$ more efficiently by machine [23, 24, 25, 26].
Though the information theory problems above usually involve more than four random variables, and the corresponding faces are of dimension higher than $2$, following the characterizations of extreme rays. i.e., $1$-dimensional faces containing matroid [7, 8, 9], and $2$-dimensional faces of $\Gamma_{3}$ [10, 11], in this series of two papers, we characterize entropy functions on the $2$-dimensional faces of $\Gamma_{4}$, which may serve as stepping stones to the general cases of this problem.
@@
In Part I, we first characterize entropy functions on extreme rays containing integer polymatroids $\hat{U}_{2,5}$ and $\hat{U}_{3,5}$ (see Subsection II-B for the definition), which cannot be find in any existing literatures. We then enumerate all $59$ types of $2$-dimensional faces of $\Gamma_{4}$ by an algorithm and completely characterize $49$ types of them. Among them, $13$ types are all-entropic faces, while $7$ are non-entropic ones [^1] By [14, 15, 16], all $2$-dimensional faces containing Vámos ray $V_{8}$ are non-entropic, and they are also all of the $2$-dimensional faces crosscut by Zhang-Yeung, or any other non-Shannon-type inequalities.. The characterization of $17$ types can be obtained by the $2$ types of non-trivial faces of $\Gamma_{3}$ characterized in [10] and [11]. Of the $22$ types of non-trivial faces, $12$ will be characterized in Part I. The primary technique we adopt to characterize these non-trivial faces is the graph-coloring in [10]. For a $3$-dimensional random vector $(X_{i},i\in N_{3})$, if $X_{1}$ and $X_{2}$ are independent, and $X_{3}$ is a function of $(X_{1},X_{2})$, $X_{1}$ and $X_{2}$ can be considered distributed on the two sets of vertices $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ of a bipartite graph with edges those $(x_{1},x_{2})$ such that $p(x_{1},x_{2})>0$, and these edges are colored by $x_{3}\in\mathcal{X}_{3}$. In our work, when the number of the random variables becomes four, we color bipartite graphs by two color systems or color a tripartite graph by one color system.
@@
The remaining $10$ cases of faces will be characterized in Part II [27] of the two-paper series. Among them, $8$ faces are fully and $2$ are partially characterized. This part has a more combinatorial flavour. All integer polymatroid with rank exceeding $1$ contained by an extreme ray will be associated a combinatorial structure. Then the entropy functions on $2$-dimensional faces containing $2$ such extreme rays breed new combinatorial-design structures, generalizing corresponding combinatorial structures in both extreme rays. These structures can also be considered as generalizations of $\rm VOA$s introduced in [7, 8].
@@ II-B Faces of a polyhedral cone
each of which determines a facet of $\Gamma_{n}$ [2, Chapter 14]. When $n=4$, it can be checked that there are $28$ elemental inequalities, or $28$ facets of $\Gamma_{4}$.
@@
where $\mathbf{r}$ is the minimal integer polymatroid in the ray, that is, an integer polymatroid such that $\mathbf{r}/t$ is not integer for any integer $t>1$. .
Therefore, in this paper, we refer to the minimal integer polymatroid $\mathbf{r}$ as the extreme ray $E_{\mathbf{r}}$ containing it unless otherwise specified. Note that some of these integer polymatroids are matroids. The $41$ extreme rays can be classified into the following $11$ types.
@@
for $\mathcal{W}^{\alpha}_{2}$ with $\alpha\subseteq N_{4}$, $|\alpha|=2$, it is called a wheel matroid with order 22, [^2] We adopt the notation and terminology in [28, Section 8.4] and it is a matroid with two parallel elements in $\alpha$, and each element in $\alpha$ and the other two elements in $N_{4}$ form a $U_{2,3}$;
@@
for $\hat{U}^{i}_{2,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{2,5}$, [^3] See [31, Theorem 1.3.6] or [32, Theorem 4] for the definition of free expansion. and its rank function is defined by
@@
for $\hat{U}^{i}_{3,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{3,5}$, and its rank function is defined by
@@
for $V^{\alpha}_{8}$ with $\alpha\subseteq N_{4}$ and $|\alpha|=2$, it is a polymatroid whose free expansion is the Vámos matroid, and its rank function is defined by
@@
It can be seen that for an extreme ray in the form $E^{\alpha}$ with $\alpha\subseteq N_{4}$, it is in a type with $\binom{4}{|\alpha|}$ extreme rays, and each extreme ray in the type can be obtained from each other by permuting the indices in $N_{4}$. For a specific extreme ray, say $U^{\{1,2\}}_{1,2}$, for simplicity, we will drop the bracket and comma of the set in the superscript, and write it as $U^{12}_{1,2}$. To facilitate the readers, the $11$ types of extreme rays are presented in Table 1 by the rank functions of their minimal integer polymatroid in the form of $15$-dimensional vectors, i.e., $\mathbf{r}=(\mathbf{r}(A),\emptyset\neq A\subseteq N_{4})$.
@@
Table 1: Extreme rays of $\Gamma_{4}$ and their rank functions
[…]
$U^{1}_{1,1}$ 11 0 0 0 1 11 11 0 0 0 11 11 11 11 11
$U^{12}_{1,2}$ 11 11 0 0 11 11 11 11 11 0 11 11 11 11 11
$U^{123}_{1,3}$ 11 11 11 0 11 11 11 11 11 11 11 11 11 11 11
$U_{1,4}$ 11 11 11 11 11 11 11 11 11 11 11 11 11 11 11
$U^{123}_{2,3}$ 11 11 11 0 22 22 11 22 11 11 22 22 22 22 22
$\mathcal{W}^{12}_{2}$ 11 11 11 11 11 22 22 22 22 22 22 22 22 22 22
$U_{2,4}$ 11 11 11 11 22 22 22 22 22 22 22 22 22 22 22
$U_{3,4}$ 11 11 11 11 22 22 22 22 22 22 33 33 33 33 33
$\hat{U}^{1}_{2,5}$ 22 11 11 11 22 22 22 22 22 22 22 22 22 22 22
$\hat{U}^{1}_{3,5}$ 22 11 11 11 33 33 33 22 22 22 33 33 33 33 33
$V^{12}_{8}$ 22 22 22 22 44 33 33 33 33 33 44 44 44 44 44
@@ II-C Entropy functions on the extreme rays of $\Gamma_{4}$
In the following, extreme rays are written in the form in (II.B.3). For each type of extreme ray, we only consider one of them. Throughout this paper, for any random vector $(X_{i}:i\in N_{4})$, each $X_{i}$ is assumed distributed on a finite set $\mathcal{X}_{i},i\in N_{4}$. For any marginal distribution of $(X_{i}:i\in N_{4})$, say $p_{X_{2}X_{3}}(x_{2},x_{3})$ with $(x_{2},x_{3})\in\mathcal{X}_{2}\times\mathcal{X}_{3}$, we usually drop its subscription and write it as $p(x_{2},x_{3})$ if there is no ambiguity. We may further simplify it as $p(x_{23})$.
@@
For $E=U^{123}_{2,3},\mathcal{W}^{14}_{2},U_{3,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log v$ for integer $v\geq 1$.
@@
([7, Proposition 2]) For $E=U_{2,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log v$ for positive integer $v\neq 2,6$.
@@
Extreme rays $E$ Theorems Figures of entropy region $E^{*}=E\cap\Gamma^{*}_{4}$ Entropy region $E^{*}$
@@
$U^{\alpha}_{2,3}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\mathcal{W}^{\alpha}_{2}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$U_{2,4}$ II-C.3 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,v>0,v\in\mathbb{Z},v\neq 2,6\}$
$U_{3,4}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{2,5}$ II-C.4 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{3,5}$ II-C.5 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
@@
It is well known that the rays containing a minimal integer polymatroid with rank-$1$ are all-entropic. It has been proven that the vectors $a\mathbf{r}$ in $U^{\alpha}_{2,3}$ are entropic if and only if $a=\log{v}$ for $k\in\mathbb{Z}$ [2, Theorem 15.2]. In the ray $U_{2,4}$, $a$ can take the values $\log{v}$ except for $\log 2$ and $\log 6$ [8, Proposition 6]. In the ray $U_{3,4}$, the region is the same as in $U^{\alpha}_{2,3}$. These three cases correspond to the Latin Square, Orthogonal Two Latin Squares, and Latin Cube, respectively. As for the ray $\hat{U}^{i}_{2,5}$ and $\hat{U}^{i}_{3,5}$, they are implied by the following theorems. For the ray $V^{\alpha}_{8}$, the fact that it disobeys the Zhang-Yeung inequality indicates that it lies outside of $\overline{\Gamma_{4}^{*}}$ and, consequently, is non-entropic.
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v>0$.
@@
For $\mathbf{h}=a\mathbf{r}$, where $\mathbf{r}$ is the rank function of $\hat{U}^{1}_{2,5}$, if we restrict it on $\{2,3,4\}$, we will see that its restriction $\mathbf{h}^{\prime}=a\mathbf{r}_{1}$, where $\mathbf{r}_{1}$ is the rank function of $U_{2,3}$, is entropic only if $a=\log v$ for integer $v>0$.
To prove the “if” part, let $X_{2}$ and $X_{3}$ be independent and uniformly distributed on $\mathbb{I}_{v}\triangleq\{0,\dots,v-1\}$, and $X_{4}=X_{2}+X_{3}\mod v$. Let $X_{1}=v\cdot X_{2}+X_{3}$. Then $X_{1}$ is uniformly distributed on $\mathbb{I}_{v^{2}}$ and it can be checked that $(X_{i},i\in N_{4})$ such constructed characterizes $\mathbf{h}=\log{v}\cdot\mathbf{r}$. ∎
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v>0$.
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{i,j,k}),$
$\qquad\qquad\qquad\quad\ \{i,j,k\}=\{2,3,4\}$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\qquad|K|=2,K\nsubseteq\{2,3,4\}.$
Assume that each $X_{i}$ is distributed on $\mathcal{X}_{i}$ and for $x_{i}\in\mathcal{X}_{i}$, $p(x_{i})>0$. For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
By Lemma 1, $X_{2},X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{2},\mathcal{X}_{3}$ and $\mathcal{X}_{4}$ and $H(X_{2})=H(X_{3})=H(X_{4})=\log v$, where $v=|\mathcal{X}_{i}|,i\in\{2,3,4\}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have $H(X_{2})=H(X_{3})=H(X_{4})=a$. Thus $a$ can only take the value $\log v$.
Now we prove that all of these polymatroids are entropic. For $a=\log v$, let $X_{2},X_{3}$ and $X_{4}$ be mutually independent and uniformly distributed on $\mathbb{I}_{v}$. Let
$X_{1}=v$ $[(X_{3}-X_{2})\mod v]$
$+$ $[(X_{4}-X_{2})\mod v].$ (II.C.23)
It remains to check the entropy function of such constructed $(X_{i},i\in N_{4})$ is $a\mathbf{r}$. Given that $X_{2}$, $X_{3}$ and $X_{4}$ are mutually independent and uniformly distributed, we have
$H(X_{2})$ $=H(X_{3})=H(X_{4})=\log v$ (II.C.24)
$H(X_{234})$ $=H(X_{2})+H(X_{3})+H(X_{4})$ (II.C.25)
$=3\log v.$ (II.C.26)
Then by (II.C.23), we obtain $H(X_{1234})=H(X_{234})$. Dividing both sides of (II.C.23) by $v$ and then rounding them down, we obtain
$\lfloor\dfrac{X_{1}}{v}\rfloor=(X_{3}-X_{2})\mod v$ (II.C.27)